A balloon contains 7.5 L of helium gas at 25°C under atmosphere pressure. If the balloon is placed in a freezer that has temperature of -4.6°C inside, to what volume will the balloon shrinks

1
Expert's answer
2022-02-07T17:11:07-0500

pV = nRT

Atmospheric pressure p = 101 325 Pa

T1 = 25 + 273 = 298 K

T2 = -4.6 + 273 = 268.4 K

R = 8.314 m3⋅Pa⋅K−1⋅mol−1

n = pV / RT = (101 325 x 0.0075) / (8.314 x 298) = 0.31 mol

V = nRT / p = (0.31 x 8.314 x 268.4) / 101 325 = 0.0068 m3 = 6.8 L